# Designing Elegant Bell Inequalities

### v1

@@ III Main result
when $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7). Saturation condition is given as $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}W_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=\tau^{n}(W_{1}^{n},W_{2}^{n},\ldots,W_{d^{2}-1}^{n})^{\intercal}$. The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after. In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$. The solution is derived in the below theorem.
@@
Coefficients in Eq.(9) also satisfies the decomposition condition. Substituting (9) into $\mathbf{r}_{a}^{n}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$. Finally, showing the saturation condition is always solvable completes the theorem. The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [34]. It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$. $\blacksquare$
@@ IV.2 Bell inequality for two-qutrit system
Substituting the derived coefficients into the expression (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
@@ V Conclusion
We presented a method to construct correlation function for which the quantum maximum is always obtained in a simple analytic form from the generalized EBI correlation. Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI. Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI cryptographic protocols based on it.
@@ Appendix B Derivation of the quantum bound
Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is to be presented here. The SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as

### v2

@@ III Main result
when $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ for all cases of $a\neq x$.
Second, we solve the saturation condition with optimal realization, $\left|\phi_{d}^{+}\right>$ and observables (5) and (7). Saturation condition is given as $(B_{y}^{n})^{*}=D_{y}^{n}=\sum_{x}f_{x,y}^{n}W_{x}^{n}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ where $\mathbf{A}_{n}:=\tau^{n}(W_{1}^{n},W_{2}^{n},\ldots,W_{d^{2}-1}^{n})^{\intercal}$. The solution $\mathbf{c}_{y}^{n}$ completely derives the coefficient matrices we after. In sum, we have to solve $(B_{y}^{n})^{*}=\mathbf{c}_{y}^{n}\cdot\mathbf{A}_{n}$ $\forall y,n$ about $\{f_{x,y}^{n}\}$ under constraint $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=0$ $\forall a\neq x$. The solution is derived in the below theorem.
@@
Coefficients in Eq.(9) also satisfies the decomposition condition. Substituting (9) into $(\mathbf{r}_{a}^{n})^{*}\cdot\mathbf{r}_{x}^{n}=\sum_{y}(f_{a,y}^{n})^{*}f_{x,y}^{n}$, one can find that it reduces to 0 for all $n$ if $a\neq x$. Finally, showing the saturation condition is always solvable completes the theorem. The elements of $\mathbf{A}_{n}$ is the unitary operator bases for traceless operator [34]. It means that the traceless operator $(B_{0}^{n})^{*}$ can always be expressed as $\mathbf{c}_{0}^{n}\cdot\mathbf{A}_{n}$ with suitable $\mathbf{c}_{0}^{n}$. $\blacksquare$
@@ IV.2 Bell inequality for two-qutrit system
Substituting the derived coefficients into (2), a Bell inequality maximally violated by generalized EBI correlation in two-qurtrit system is derived as,
@@ V Conclusion
We presented a method to construct correlation function for which the quantum maximum is always obtained in a simple analytic form from the generalized EBI correlation. Bell inequality for two-qutrit system derived in our method shows larger violation compared to known Bell inequalities of similar classes, including original EBI. Our work lays the foundation for the future studies for self-testing MUBs and SICs in high dimensional system and DI quantum cryptographic protocols based on it.
@@ Appendix B Derivation of the quantum bound
Detailed version of the derivation of quantum upper bound, $\tilde{Q}$, is presented in this section. As shown in the main text, the SOS decomposition $\sum_{n,y}(P_{y}^{n})^{\dagger}P_{y}^{n}\geq 0$ can be modified as